# Epidemic threshold and localization of the SIS model on directed complex networks

### v1

@@ Abstract
We study the susceptible-infected-susceptible (SIS) model on directed complex networks within the quenched mean-field approximation. Combining results from random matrix theory with an analytic approach to the distribution of fixed-point infection probabilities, we derive the phase diagram and show that the model exhibits a nonequilibrium phase transition between the absorbing and endemic phases for $c\geq\lambda^{-1}$, where $c$ is the mean degree and $\lambda$ the average infection rate. Interestingly, the critical line is independent of the degree distribution but is highly sensitive to the form of the infection-rate distribution. We further show that the inverse participation ratio of infection probabilities diverges near the epidemic threshold, indicating that the disease may become localized on a small fraction of nodes. These results provide a comprehensive picture of how network heterogeneities shape epidemic spreading on directed contact networks.
@@ I Introduction
where $\Lambda_{1}(N)>0$ denotes the leading eigenvalue (spectral radius) of the network adjacency matrix Newman (2010). Beyond its role in QMF theory, Eq. (1) provides a rigorous lower bound for the exact epidemic threshold on any given network Van Mieghem and van de Bovenkamp (2013).
@@
In contrast to undirected networks, the SIS model on directed contact networks, where the infection rates $\lambda_{ij}$ are unidirectional  [^1] In a directed contact network, if node $j$ infects node $i$ with rate $\lambda_{ij}$, the opposite rate $\lambda_{ji}$ is zero., remains poorly understood. Spreading models on directed networks provide a natural framework for studying the transmission of computer viruses through email networks Newman et al. (2002); Ebel et al. (2002), the diffusion of information in social networks Krueger et al. (2024), and the transmission of diseases between patients and health care workers Meyers et al. (2006). Only a few works have examined how directionality impacts the SIS model Li et al. (2013); Kwon and Kim (2013). Numerical studies on finite networks with both directed and bidirected edges have shown that $\lambda_{\rm c}(N)$ increases with the fraction of directed links Li et al. (2013). In addition, the SIS model has been studied on directed networks using heterogeneous mean-field theory Kwon and Kim (2013), which does not fully capture the underlying network structure. As a result, even basic problems, such as how the directed network structure affects the epidemic threshold, remain unresolved.
@@ II The SIS model on directed networks
$\frac{d\rho_{i}}{dt}=-\rho_{i}(t)+\left[1-\rho_{i}(t)\right]\sum_{j=1(\neq i)}^{N}A_{ij}\rho_{j}(t),$ (2)
where the elements $\{A_{ij}\}_{i,j=1,\dots,N}$ of the $N\times N$ weighted adjacency matrix $\bm{A}$ have the form $A_{ij}=C_{ij}\lambda_{ij}$, with $C_{ij}\in\{0,1\}$ and $\lambda_{ij}>0$. The binary variables $\{C_{ij}\}_{i,j=1,\dots,N}$ specify the structure of the contact network: if $C_{ij}=1$, there is a directed link $j\rightarrow i$ from node $j$ to $i$, whereas $C_{ij}=0$ otherwise. The indegree $K_{i}$ (outdegree $L_{i}$) of node $i$ is given by $K_{i}=\sum_{j=1}^{N}C_{ij}$ ($L_{i}=\sum_{j=1}^{N}C_{ji}$). We consider directed random networks generated from the configuration model Newman et al. (2001); Fosdick et al. (2018), where the degree sequences $K_{1},\dots,K_{N}$ and $L_{1},\dots,L_{N}$ are independent and identically distributed random variables drawn from $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$, with $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ denoting, respectively, the indegree and outdegree distributions. The average degree $c$ is determined from
@@
The coefficient $\lambda_{ij}>0$ denotes the infection rate that node $j$ infects $i$. We assume that $\{\lambda_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables sampled from a distribution $P_{\lambda}(x=\lambda_{ij})$ with mean $\lambda$ and variance $\sigma^{2}$. We will present results for two choices of $P_{\lambda}(x)$. The first one is the $\Gamma$-distribution
@@ III The cavity approach for the distribution of infection probabilities
In this section, we present a first-principles derivation of Eq. (11) for the distribution of the fixed-point infection probabilities $\rho_{1},\dots,\rho_{N}$ in the limit $N\rightarrow\infty$. This section can be skipped by readers primarily interested in the results.
@@
$F_{i}(\bm{\rho})=-\rho_{i}+(1-\rho_{i})\sum_{j=1}^{N}A_{ij}\rho_{j}=0,$ (12)
where $i=1,\dots,N$. Assuming that, for a single realization of $\bm{A}$, Eq. (12) admits a unique solution, the joint probability density of $\bm{\rho}$ can be formally written as Kurchan (1991)
$\mathcal{P}_{N}(\bm{\rho})=\dfrac{\prod_{i=1}^{N}\delta\left[F_{i}(\bm{\rho})\right]}{\int\limits_{0}^{1}\left(\prod_{i=1}^{N}d\rho_{i}\right)\prod_{i=1}^{N}\delta\left[F_{i}(\bm{\rho})\right]}.$ (13)
Using the Fourier transform of the Dirac-$\delta$, we write $\mathcal{P}_{N}(\bm{\rho})$ as
@@
where
$\gamma_{N}(\bm{\rho},\hat{\bm{\rho}})=\frac{\exp{\left[H_{N}(\bm{\rho},\hat{\bm{\rho}})\right]}}{\int\limits_{-\infty}^{\infty}\left(\prod_{i=1}^{N}d\hat{\rho}_{i}\right)\int\limits_{0}^{1}\left(\prod_{i=1}^{N}d\rho_{i}\right)\exp{\left[H_{N}(\bm{\rho},\hat{\bm{\rho}})\right]}}$, (15)
with
$H_{N}(\bm{\rho},\hat{\bm{\rho}})=-i\sum_{j=1}^{N}\hat{\rho}_{j}\rho_{j}+i\sum_{lj=1}^{N}A_{lj}\hat{\rho}_{l}(1-\rho_{l})\rho_{j}.$ (16)
Our purpose is to compute the local marginals $\{\mathcal{P}_{N,i}(\rho_{i})\}_{i=1}^{N}$ on the network nodes by using the cavity method Metz et al. (2010).
From Eq. (14), the marginal $\mathcal{P}_{N,i}(\rho_{i})$ is given by
$\mathcal{P}_{N,i}(\rho_{i})=\int\limits_{-\infty}^{\infty}d\hat{\rho}_{i}\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}).$ (17)
@@
$H_{N}(\bm{\rho},\hat{\bm{\rho}})=-i\hat{\rho}_{i}\rho_{i}+i\rho_{i}\sum_{j\in\partial_{i}^{\rm out}}\lambda_{ji}\hat{\rho}_{j}(1-\rho_{j})$
[…]
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}\int\limits_{-\infty}^{\infty}\Bigg{(}\prod_{j\in\partial_{i}}d\hat{\rho}_{j}\Bigg{)}\int\limits_{0}^{1}\Bigg{(}\prod_{j\in\partial_{i}}d\rho_{j}\Bigg{)}$
@@
$\times\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}),$ (19)
@@
where $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}})$ is defined on the neighbourhood $\partial_{i}$ within the cavity graph $\mathcal{G}_{N-1}^{(i)}$. We have omitted the normalization constant in Eq. (19), as this quantity can be fixed at the end of the calculation.
@@
At this point, we invoke the main assumption of the cavity method on sparse random graphs Mézard and Parisi (2001); Metz et al. (2010). Since networks generated from the configuration model become locally tree-like for $N\gg 1$, the function $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}})$ factorizes as follows
$\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}})=\prod_{j\in\partial_{i}}\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}),$ (20)
leading to the following expression for $\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})$,
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}$
@@
$\times\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j})$
$\times\prod_{j\in\partial_{i}^{\rm out}}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{ji}\rho_{i}}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}).$ (21)
In the terminology of the message-passing algorithm Mézard and Montanari (2009), $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j})$ denotes the message that propagates from node $j$ to $i$ along the directed edge $j\rightarrow i$. According to Eq. (21), the local marginal $\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})$ is determined by messages arriving at node $i$ through both ingoing and outgoing edges. However, the infection probability $\rho_{i}$ is solely determined by the in-neighbourhood $\partial_{i}^{\rm in}$, so it is sensible to assume that the messages from $\partial_{i}^{\rm out}$ contribute with a constant in Eq. (21). While we could not demonstrate this fact in full generality, making this assumption leads to results that are fully consistent with the exact solution for the dynamics of the SIS model Metz (2025).
Thus, by integrating Eq. (21) over $\hat{\rho}_{i}$ and using Eq. (17), we find
@@
where $Z_{N,i}$ ensures that $\mathcal{P}_{N,i}(\rho)$ is normalized. Hence, for large $N$, $\mathcal{P}_{N,i}(\rho)$ follows from the local marginals $\{\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\}_{j\in\partial_{i}^{\rm in}}$ on $\mathcal{G}_{N-1}^{(i)}$. To determine the cavity marginals $\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})$, we follow exactly the same reasoning as explained above, which leads to the cavity equations
@@
with $l\in\partial_{i}^{\rm out}$. The constant $Z_{N-1,i}^{(l)}$ normalizes $\mathcal{P}_{N-1,i}^{(l)}(\rho)$.
@@
The constants $Z$ and $Z_{\rm edg}$ denote the corresponding normalization factors. Equations (26) and (27) are valid for networks with an arbitrary joint distribution $p_{k\ell}$ of indegrees and outdegrees. Setting $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$, we recover Eq. (11).

### v2

@@ Abstract
We study the susceptible-infected-susceptible (SIS) model on directed complex networks within the quenched mean-field approximation. Combining results from random matrix theory with an analytic approach to the distribution of fixed-point infection probabilities, we derive the phase diagram and show that the model exhibits a nonequilibrium phase transition between the absorbing and endemic phases for $c\geq\lambda^{-1}$, where $c$ is the mean degree and $\lambda$ the average infection rate. Interestingly, the critical line is independent of the degree distribution but is highly sensitive to the form of the infection-rate distribution. We further show that the inverse participation ratio of infection probabilities diverges near the epidemic threshold, indicating that the disease may become localized on a small fraction of nodes. These results provide a systematic characterization of how network heterogeneities shape epidemic spreading on directed contact networks within the quenched mean-field approximation.
@@ I Introduction
where $\Lambda_{1}(N)>0$ denotes the leading eigenvalue (spectral radius) of the network adjacency matrix Newman (2010). Beyond its role in QMF theory, Eq. (1) provides a rigorous lower bound for the exact epidemic threshold on any given network Van Mieghem and van de Bovenkamp (2013). The QMF approximation can be improved by incorporating pairwise dynamical correlations, leading to the so-called pair quenched mean-field theory Mata and Ferreira (2013), which yields more accurate predictions for the epidemic threshold compared to the individual-based QMF theory.
@@
In contrast to undirected networks, the SIS model on directed contact networks, where the coupling strengths or infection rates $\lambda_{ij}$ are unidirectional  [^1] In a directed contact network, if node $j$ infects node $i$ with rate $\lambda_{ij}$, the opposite rate $\lambda_{ji}$ is zero., remains poorly understood. Spreading models on directed networks provide a natural framework for studying the transmission of computer viruses through email networks Newman et al. (2002); Ebel et al. (2002), the diffusion of information in social networks Krueger et al. (2024), and the transmission of diseases between patients and health care workers Meyers et al. (2006). Only a few works have examined how directionality impacts the SIS model Li et al. (2013); Kwon and Kim (2013). Numerical studies on finite networks with both directed and bidirected edges have shown that $\lambda_{\rm c}(N)$ increases with the fraction of directed links Li et al. (2013). In addition, the SIS model has been studied on directed networks using heterogeneous mean-field theory Kwon and Kim (2013), which does not fully capture the underlying network structure. As a result, even basic problems, such as how the directed network structure affects the epidemic threshold, remain unresolved.
@@ II The SIS model on directed networks
$\frac{d\rho_{i}}{dt}=-\rho_{i}(t)+\left[1-\rho_{i}(t)\right]\sum_{j=1}^{N}A_{ij}\rho_{j}(t),$ (2)
where the elements $\{A_{ij}\}_{i,j=1,\dots,N}$ of the $N\times N$ weighted adjacency matrix $\bm{A}$ have the form $A_{ij}=C_{ij}\lambda_{ij}$, with $C_{ij}\in\{0,1\}$ and $\lambda_{ij}>0$. The binary variables $\{C_{ij}\}_{i,j=1,\dots,N}$ specify the structure of the contact network: if $C_{ij}=1$, there is a directed link $j\rightarrow i$ from node $j$ to $i$, whereas $C_{ij}=0$ otherwise. The diagonal elements $\{C_{ii}\}_{i=1,\dots,N}$ are zero. The indegree $K_{i}$ (outdegree $L_{i}$) of node $i$ is given by $K_{i}=\sum_{j=1}^{N}C_{ij}$ ($L_{i}=\sum_{j=1}^{N}C_{ji}$). We consider directed random networks generated from the configuration model Newman et al. (2001); Fosdick et al. (2018), where the degree sequences $K_{1},\dots,K_{N}$ and $L_{1},\dots,L_{N}$ are independent and identically distributed random variables drawn from $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$, with $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ denoting, respectively, the indegree and outdegree distributions. The degree sequences fulfill the constraint $\sum_{i=1}^{N}K_{i}=\sum_{i=1}^{N}L_{i}$, and the average degree $c$ reads
@@
The coefficient $\lambda_{ij}>0$ denotes the directed coupling strength $j\rightarrow i$ or the pairwise infection rate at which node $j$ infects $i$. We assume that $\{\lambda_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables sampled from a distribution $P_{\lambda}(x=\lambda_{ij})$ with mean $\lambda$ and variance $\sigma^{2}$. We will present results for two choices of $P_{\lambda}(x)$. The first one is the $\Gamma$-distribution
@@ III The cavity approach for the distribution of infection probabilities
In this section, we present a first-principles derivation of Eq. (11) using the cavity method from spin-glass theory. The dynamic cavity method has been developed in the context of epidemic spreading to estimate the probability $\rho_{i}(t)$ that node $i$ is infected at time $t$ Lokhov et al. (2014, 2015); Shrestha et al. (2015); Ortega et al. (2022), leading to important applications such as the inference of the origin of an epidemic outbreak Lokhov et al. (2014); Altarelli et al. (2014a) and the design of optimal immunization strategies Altarelli et al. (2014b). Here, instead of following a dynamical approach, we start from the fixed-point equations for the infection probabilities $\rho_{1},\dots,\rho_{N}$ in the QMF approximation and compute their empirical distribution using the cavity method. We note that this section can be skipped by readers primarily interested in the results obtained by solving Eq. (11).
@@
$F_{N,i}(\bm{\rho}|\bm{h})=-\rho_{i}+(1-\rho_{i})\sum_{j=1}^{N}A_{ij}\rho_{j}+h_{i}=0,$ (12)
where $i=1,\dots,N$. We have added auxiliary fields $\bm{h}=(h_{1},\dots,h_{N})^{T}$ to the fixed-point equations, which are useful to identify the physical meaning of certain quantities. These fields will be set to zero at the end of the calculation. The joint probability density of $\bm{\rho}$ can be formally written as Kurchan (1991)
$\mathcal{P}_{N}(\bm{\rho}|\bm{h})=\dfrac{\prod_{i=1}^{N}\delta\left[F_{N,i}(\bm{\rho}|\bm{h})\right]}{\int\limits_{0}^{1}\left(\prod_{i=1}^{N}d\rho_{i}\right)\prod_{i=1}^{N}\delta\left[F_{N,i}(\bm{\rho}|\bm{h})\right]}.$ (13)
Assuming that Eqs. (12) admit a unique stable solution, Eq. (13) concentrates as $N\rightarrow\infty$ on the joint distribution of infection probabilities characterizing this stable state. Starting from Eq. (13), our purpose is to compute the local marginals $\{\mathcal{P}_{N,i}(\rho_{i})\}_{i=1}^{N}$ on the network nodes by using the cavity method Metz et al. (2010). Using the Fourier transform of the Dirac-$\delta$, we write $\mathcal{P}_{N}(\bm{\rho}|\bm{h})$ as
$\mathcal{P}_{N}(\bm{\rho}|\bm{h})=\int_{-\infty}^{\infty}\left(\prod_{i=1}^{N}d\hat{\rho}_{i}\right)\gamma_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h}),$ (14)
where
$\gamma_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=\left[Z_{N}(\bm{h})\right]^{-1}\exp{\left[H_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})\right]},$ (15)
with
$H_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=i\sum_{j=1}^{N}h_{j}\hat{\rho}_{j}-i\sum_{j=1}^{N}\hat{\rho}_{j}\rho_{j}+i\sum_{l,j=1}^{N}A_{lj}\hat{\rho}_{l}(1-\rho_{l})\rho_{j}.$ (16)
@@
From Eq. (14), the marginal $\mathcal{P}_{N,i}(\rho_{i}|h_{i})$ is given by
$\mathcal{P}_{N,i}(\rho_{i}|h_{i})=\int\limits_{-\infty}^{\infty}d\hat{\rho}_{i}\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}|h_{i}).$ (17)
@@
$H_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=ih_{i}\hat{\rho}_{i}-i\hat{\rho}_{i}\rho_{i}+i\rho_{i}\sum_{j\in\partial_{i}^{\rm out}}\lambda_{ji}\hat{\rho}_{j}(1-\rho_{j})$
$+i\hat{\rho}_{i}(1-\rho_{i})\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}+H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h}).$ (18)
The function $H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})$ is defined on the cavity graph $\mathcal{G}_{N-1}^{(i)}$, which is obtained from the original graph $\mathcal{G}_{N}$ by removing node $i$, its adjacent edges, and the local field $h_{i}$. The symbol $\partial_{i}^{\rm out}$ denotes the set of nodes that receive a directed link from $i$ (the out-neighborhood of $i$), while $\partial_{i}^{\rm in}$ represents the set of nodes that have a directed link pointing to $i$ (the in-neighborhood of $i$). The full neighborhood of $i$ is defined as $\partial_{i}=\partial_{i}^{\rm out}\cup\partial_{i}^{\rm in}$. By substituting Eq. (18) in Eq. (15) and integrating the resulting expression over all variables except $(\rho_{i},\hat{\rho}_{i})$, we find
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}|h_{i})\sim e^{i\hat{\rho}_{i}h_{i}-i\hat{\rho}_{i}\rho_{i}}\int\limits_{-\infty}^{\infty}\Bigg(\prod_{j\in\partial_{i}}d\hat{\rho}_{j}\Bigg)\int\limits_{0}^{1}\Bigg(\prod_{j\in\partial_{i}}d\rho_{j}\Bigg)$
@@
$\times\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}}),$ (19)
@@
where $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}})$ is the joint distribution defined on the neighbourhood $\partial_{i}$ within the cavity graph $\mathcal{G}_{N-1}^{(i)}$. We have omitted the normalization constant in Eq. (19), as this quantity can be fixed at the end of the calculation.
@@
At this point, we invoke the main assumption of the cavity method on sparse random graphs Mézard and Parisi (2001); Metz et al. (2010). Since networks generated from the configuration model become locally tree-like for $N\gg 1$, the function $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}})$ factorizes as follows
$\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}})=\prod_{j\in\partial_{i}}\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j}).$ (20)
Inserting the above expression in Eq. (19) and using Eq. (17), we find
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}|h_{i})\sim e^{i\hat{\rho}_{i}h_{i}-i\hat{\rho}_{i}\rho_{i}}$
$\times\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\mathcal{P}_{N-1,j}^{(i)}(\rho_{j}|h_{j})$
$\times\prod_{j\in\partial_{i}^{\rm out}}\int\limits_{0}^{1}d\rho_{j}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{ji}\rho_{i}}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j}).$ (21)
In the terminology of the message-passing algorithm Mézard and Montanari (2009), $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$ denotes the message that propagates from node $j$ to $i$ along the directed edge $j\rightarrow i$. Since a node $i$ can be infected only by its in-neighbourhood $\partial_{i}^{\rm in}$, the term containing the messages $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$ originated in the out-neighbourhood $j\in\partial_{i}^{\rm out}$ of $i$ should be irrelevant and contribute with a constant in Eq. (21).
Let us now demonstrate this property. By repeating the same reasoning that led to Eq. (21), we can derive the cavity equations for $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$. For $i\in\partial_{j}^{\rm in}$, we obtain
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})\sim e^{i\hat{\rho}_{j}h_{j}-i\hat{\rho}_{j}\rho_{j}}$
$\times\prod_{l\in\partial_{j}^{\rm in}\setminus i}\,\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{jl}\rho_{l}}\,\mathcal{P}_{N-1,l}^{(j)}(\rho_{l}|h_{l})$
$\times\prod_{l\in\partial_{j}^{\rm out}}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{l}\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{l}(1-\rho_{l})\lambda_{lj}\rho_{j}}\,\gamma_{N-1,l}^{(j)}(\rho_{l},\hat{\rho}_{l}|h_{l}),$ (22)
while for $i\in\partial_{j}^{\rm out}$ we get
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})\sim e^{i\hat{\rho}_{j}h_{j}-i\hat{\rho}_{j}\rho_{j}}$
$\times\prod_{l\in\partial_{j}^{\rm in}}\,\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{jl}\rho_{l}}\,\mathcal{P}_{N-1,l}^{(j)}(\rho_{l}|h_{l})$
@@
$\times\prod_{l\in\partial_{j}^{\rm out}\setminus i}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{l}\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{l}(1-\rho_{l})\lambda_{lj}\rho_{j}}\,\gamma_{N-1,l}^{(j)}(\rho_{l},\hat{\rho}_{l}|h_{l}),$ (23)
where $\partial_{j}^{\rm in}\setminus i$ ($\partial_{j}^{\rm out}\setminus i$) represents the set of nodes in the in-neighbourhood (out-neighbourhood) of $j$, except for node $i$. Let us focus on $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$ with $i\in\partial_{j}^{\rm in}$, which is precisely the quantity appearing in the right hand side of Eq. (21). By setting all auxiliary fields to zero in Eq. (22), except for $h_{j}$, we obtain the interesting identity
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})=e^{i\hat{\rho}_{j}h_{j}}\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}),$ (24)
with $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j})\equiv\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|0)$. Hence, setting all external fields to zero in Eq. (21) and using the above identity, we find
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})$
$\times\prod_{j\in\partial_{i}^{\rm out}}\int\limits_{0}^{1}d\rho_{j}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|(1-\rho_{j})\lambda_{ji}\rho_{i}).$ (25)
The final step of this calculation is to understand the physical meaning of $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|(1-\rho_{j})\lambda_{ji}\rho_{i})$. This can be achieved by extracting the $j$-th term from the function $H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})$ on the cavity graph $\mathcal{G}_{N-1}^{(i)}$,
$H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=ih_{j}\hat{\rho}_{j}-i\hat{\rho}_{j}\rho_{j}+i\rho_{j}\sum_{l\in\partial_{j}^{\rm out}}\lambda_{lj}\hat{\rho}_{l}(1-\rho_{l})$
$+i\hat{\rho}_{j}(1-\rho_{j})\sum_{l\in\partial_{j}^{\rm in}\setminus i}\lambda_{jl}\rho_{l}+H_{N-2}^{(i,j)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h}),$ (26)
where $i\in\partial_{j}^{\rm in}$. If we choose $h_{j}=(1-\rho_{j})\lambda_{ji}\rho_{i}$ in the above equation, we restore the neighbourhood $\partial_{j}$ of $j$ as in the original graph. In other words, by setting the external field to $h_{j}=(1-\rho_{j})\lambda_{ji}\rho_{i}$, we add back the interaction term between $j$ and $i$, which has been deleted through the removal of $i$. Hence, for $N\gg 1$, the following relation holds
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|(1-\rho_{j})\lambda_{ji}\rho_{i})=\gamma_{N,j}(\rho_{j},\hat{\rho}_{j}).$ (27)
Inserting the above relation in Eq. (25) and using the normalization of $\gamma_{N}(\rho_{j},\hat{\rho}_{j})$, we obtain
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\mathcal{P}_{N-1,j}^{(i)}(\rho_{j}),$ (28)
as expected.
Finally, by integrating Eq. (28) over $\hat{\rho}_{i}$, we find
@@
where $Z_{N,i}$ ensures that $\mathcal{P}_{N,i}(\rho)$ is normalized. Hence, for large $N$, $\mathcal{P}_{N,i}(\rho)$ follows from the local marginals $\{\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\}_{j\in\partial_{i}^{\rm in}}$ on $\mathcal{G}_{N-1}^{(i)}$. To determine the cavity marginals $\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})$, we start from Eq. (23) and follow exactly the same reasoning as explained above, which leads to the cavity equations
@@
with $l\in\partial_{i}^{\rm out}$. The constant $Z_{N-1,i}^{(l)}$ normalizes $\mathcal{P}_{N-1,i}^{(l)}(\rho)$. The local marginals $\mathcal{P}_{N,i}(\rho)$ and $\mathcal{P}_{N-1,i}^{(l)}(\rho)$ fulfill the same equation on a single graph instance because the quantity $\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}$ is solely determined by the in-neighbourhood $\partial_{i}^{\rm in}$, while $\mathcal{P}_{N-1,i}^{(l)}(\rho)$ is obtained by removing a node $l$ from the out-neighbourhood. This is different, for instance, from the cavity approach in undirected networks, where there is no distinction between in-neighbourhood and out-neighbourhood, and the removal of a node influences the states of all its neighbours.
@@
The constants $Z$ and $Z_{\rm edg}$ denote the corresponding normalization factors. Equations (33) and (34) are valid for networks with an arbitrary joint distribution $p_{k\ell}$ of indegrees and outdegrees. Substituting $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$ in Eq. (34) and summing over $l$, we conclude that $\mathcal{P}(\rho)=\mathcal{P}_{\rm edg}(\rho)$, recovering Eq. (11).